You read about a full time summer position in sales at the Furniture barn. The job pays$260 per week plus 20% commission on sales over $1000. Explain how you would calculate the total amount earned each week.

Answer by JulietG(1812)

Base + Commission = Pay
260 + (.2 * $>$1k) = pay
Let's say you sold $500 in your first week.
Your pay would be 260 + 0 = $260
You do better in your second week, selling $1250
Your pay would be 260 + (.2 * 250) = $310
In your third week, you have the hang of it and sell $3000 worth
Your pay would be 260 + (.2 * 2000) = $660
Week four there's a huge clearance sale where you sell a whopping $5000
Your pay would be $260 + (.2 * 4000) = $1060